# Unbounded-width CSPs are Untestable in a Sublinear Number of Queries

### v1

@@ 2.3 Irredundant Relations and Galois Duality
A fundamental result of Geiger [Gei68] shows that if the notion of “relations generated by $\Gamma$” in Definition 2.1 is suitably relaxed, then the resulting set $\langle\Gamma\rangle^{*}$ actually coincides with $\langle\Gamma\rangle$. This result captures a deep “duality” between algebraic and relational structures (an instance of the more general concept of a Galois connection), which we do not elaborate on here. Instead, following Dalmau and Krokhin [DK13], we state a variant of Geiger’s theorem asserting that any irredundant relation in $\langle\Gamma\rangle$ is contained in $\langle\Gamma\rangle^{*}$ (without relaxing Definition 2.1).
@@
Theorem 2.14 ([Gei68]).
Let $(D,\Gamma)$ be a relational structure and let $R:D^{k}\rightarrow\{0,1\}$ be an irredundant relation on $D$. If every polymorphism of $(D,\Gamma)$ preserves $R$, then $R\in\langle\Gamma\rangle^{*}$.
For the sake of completeness, a proof of Theorem 2.14 is given in Appendix B.
@@
For any relational structure $(D,\Gamma)$, we have $End_{\Gamma}\in\langle\Gamma\rangle^{*}$.
@@ 3 Main Lemma from Universal Algebra
The statement of Theorem 3.2 is primarily derived from [BK14, Conjecture 4.3], with additional ingredients drawn from other works, such as [Val09]. Further details concerning Theorem 3.2 will be provided in Appendix D.
@@
Let $(D,\Gamma)$ be a relational structure of unbounded width. Then there exist a subset $D^{\prime}\subseteq D$, an Abelian group $\mathsf{G}$, and a surjective map $\varphi\colon D^{\prime}\to\mathsf{G}$ such that any ternary relation $R\colon D^{3}\to\{0,1\}$ of the form
@@ 4 Hardness of Linear Equations
Theorem 4.1 ([BOT02]).
For any Abelian group $\mathsf{G}$ and any $\varepsilon>0$, there exists a constant $d>0$ such that $\mathsf{MaxCSP}(\textsc{3Sum}_{\mathsf{G}})[1,\frac{1}{|\mathsf{G}|}+\varepsilon]$ requires $\Omega(n)$ queries in the $\mathrm{BD}^{*}(d,\frac{d}{3},3n)$ model. [^11] The paper [BOT02] only proved Theorem 4.1 in the case $\mathsf{G}=\mathbb{Z}/2\mathbb{Z}$, but the proof can easily be adapted to work for general Abelian groups.
However, unlike [DK13], which applies the result of [Hås01] as a black box to prove Theorem 1.9, we cannot directly use Theorem 4.1 due to the subtleties of our bounded-degree query model. In this section, we unpack the proof of Theorem 4.1 and extract the components that can be explicitly leveraged in proving Theorem 1.8. For the sake of convenience, several aspects of the proof in [BOT02] are slightly modified in the presentation below.
@@
It is not hard to deduce Theorem 4.1 from Proposition 4.4 and Lemmas 4.5 and 4.7. For the sake of completeness, proofs of Lemmas 4.5 and 4.7 will be sketched in Appendix C.
@@ 5.1 The Expander Gadget
Recall that our universal-algebraic lemma (Lemma 3.4) guarantees the simulation of “linear equation constraints” only by $\Gamma\cup\textsc{Const}_{D}$, rather than by $\Gamma$ alone. Consequently, we must first find a way to simulate the constant relations using only the relations in $\Gamma$. Note that constant relations may not be generated by $\Gamma$ in the sense of Definition 2.1, so different techniques are required. Interestingly, the original work of [BOT02] encountered the same difficulty when proving the hardness of testing graph 3-colorability. [^12] We now view this need for constant relations in [BOT02] as a special case of the general phenomenon described in Lemma 3.4, since 3-colorability is an example of an unbounded-width CSP. To address this issue, [BOT02] uses an “expander gadget” to fix the assignments of certain variables — a technique that also plays a crucial role in our proof.
In contrast to [BOT02], where ordinary expander graphs suffice, our more general setting requires the use of random regular hypergraphs, defined formally below.
@@
A proof of Lemma 5.4 is given in Appendix C.
@@ 5.2 Construction in the Core Case
Recall that the expander gadget constructed in Section 5.1 relies on the availability of a sub-unique relation. However, some CSP templates cannot generate any sub-unique relation (in the sense of Definition 2.1). Fortunately, when the template $(D,\Gamma)$ is a core (see Definition 2.4), the “Endomorphism” relation provides a natural sub-unique relation that can be generated from $\Gamma$, as established in Proposition 2.16.141414The idea of using the Endomorphism relation is due to [DK13].
In this subsection, we construct the YES and NO distributions for Theorem 1.8 in the case the template is a core. Throughout this subsection, we fix a relational structure $(D,\Gamma)$, and assume that it is a core and has unbounded width. In particular, this implies $|D|\geqslant 2$.
@@ 5.4 Conclusion in the General Case
We now turn to the case where $(D,\Gamma)$ is not a core. By Proposition 2.6, $(D,\Gamma)$ is homomorphically equivalent to a core $(D^{\prime},\Gamma^{\prime})$. We claim that $\mathsf{MaxCSP}(\Gamma^{\prime})[1,1-\varepsilon]$ is essentially equivalent to $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$.
@@ 2   Initialize a set of partial assignments $S_{U}\leftarrow D^{U}$
Let $(D,\Gamma)$ be a relational structure of unbounded width. Then $(D,\Gamma\cup\textsc{Const}_{D})$ also has unbounded width.
@@
This is simply because any $\mathrm{CSP}(\Gamma)$ instance is also a $\mathrm{CSP}(\Gamma\cup\textsc{Const}_{D})$ instance. If Algorithm 1 makes error on some instance $I\in\mathrm{CSP}(\Gamma)$, then it also makes error on the same instance $I\in\mathrm{CSP}(\Gamma\cup\textsc{Const}_{D})$. ∎
@@
A decidable characterization of bounded-width CSP templates was given by Barto [Bar14, Corollary 8.5]. Using Barto’s criterion, it is easy to verify that for any $k,\ell\geqslant 2$ with $(k,\ell)\neq(2,2)$, the CSP template expressing $k$-colorability of $\ell$-uniform hypergraphs does not have bounded width.
Appendix B Galois Duality Proof
@@
Appendix C Random Regular Hypergraphs
In this section, we outline the proofs of Lemmas 4.5, 4.7 and 5.4, all of which revolve around random regular hypergraphs. The first two lemmas, Lemmas 4.5 and 4.7, are only minor adaptations of results from [BOT02], and our proofs closely follow those in that work. In contrast, Lemma 5.4 is new to this paper, though its proof uses arguments that are largely analogous to those used in the proof of Lemma 4.5.
C.1 Random Regular Hypergraphs are Locally Sparse
@@
Recall from Definition 4.2 that $\mathcal{M}_{[3],n}$ denotes the set of all $3$-partite perfect matchings on the vertex set $n]\times[3]$. For any matching $M\in\mathcal{M}_{[3],n}$ and any vertex subset $U\subseteq[n]\times[3]$, we let $M[U]$ denote the set of hyperedges $(v_{1},v_{2},v_{3})\in M$ such that $v_{1},v_{2},v_{3}\in U$. The following lemma is a slightly modified version of [BOT02, Lemma 6].
Lemma C.1 ([BOT02]).
@@
$\sum_{i=1}^{d}\left|M^{(i)}[U]\right|<\frac{2}{3}|U|,\quad\text{for all }U\subseteq[n]\times[3]\text{ with }|U|\leqslant\delta n$ (C.1)
@@
As stated in [BOT02], the conclusion of Lemma C.1 still holds if the constant $2/3$ in (C.1) is replaced by any constant greater than $1/2$. The reason for choosing $2/3$ in (C.1) is that it is the largest constant for which the following observation remains true.
Proposition C.2.
Suppose $M^{(1)},\dots,M^{(d)}$ are perfect matchings in $\mathcal{M}_{[3],n}$ that satisfy (C.1). Then, for any subset $U\subseteq[n]\times[3]$ with $|U|\leqslant\delta n$, the following hold:
@@
The first item holds because no hyperedge can appear in more than one of the $M^{(i)}$’s; otherwise, there would exist a set $U$ of three vertices such that $\sum_{i=1}^{d}\left|M^{(i)}[U]\right|\geqslant 2$, violating condition (C.1).
We now prove item (2). Let $U_{0}\subseteq U$ denote the set of vertices incident to at least one hyperedge in the union $E:=\bigcup_{i=1}^{d}M^{(i)}$. Applying condition (C.1) to $U_{0}$ implies that a uniformly random vertex from $U_{0}$ is, in expectation, incident to fewer than two hyperedges in $E$. Hence, there exists at least one vertex in $U_{0}$ that is incident to exactly one hyperedge in $E$.
@@
We now use Lemma C.1 and Proposition C.2 to prove Lemma 4.7.
@@
Recall from Definition 4.3 that for any sample $(I_{\mathrm{yes}},I_{\mathrm{no}})\sim\mathcal{D}_{\mathrm{pair}}^{\mathsf{G},n,d}$, the same collection of $d$ perfect matchings $M^{(1)},\dots,M^{(d)}\in\mathcal{M}_{[3],n}$ underlies both $I_{\mathrm{yes}}$ and $I_{\mathrm{no}}$. Applying Lemma C.1, we obtain a constant $\delta\in(0,1)$, and let $\mathcal{E}$ denote the event that the condition (C.1) holds for $M^{(1)},\dots,M^{(d)}$.
We now condition on a choice of $M^{(1)},\dots,M^{(d)}$ satisfying (C.1). Recall that for each hyperedge $\mathbf{v}=(v_{1},v_{2},v_{3})\in\bigcup_{i=1}^{d}M^{(i)}$, the instance $I_{\mathrm{yes}}$ contains a constraint $(\mathbf{v},S^{\mathsf{G}}_{b})$ and $I_{\mathrm{no}}$ contains a constraint $(\mathbf{v},S^{\mathsf{G}}_{b^{\prime}})$, for possibly distinct “right-hand side” coefficients $b,b^{\prime}\in\mathsf{G}$. The right-hand sides of the equations in $I_{\mathrm{no}}$ are chosen independently, whereas those in $I_{\mathrm{yes}}$ are correlated through an underlying random assignment $\tau:[n]\times[3]\to\mathsf{G}$.
However, it follows from Proposition C.2 that for any variable set $U\subseteq[n]\times[3]$ of size at most $\delta n$, the equations placed by $I_{\mathrm{yes}}$ within $U$ have right-hand sides that remain jointly independent. Consequently, conditioned on $M^{(1)},\dots,M^{(d)}$, the restricted random instances $\mathrm{Sub}_{U}[I_{\mathrm{yes}}]$ and $\mathrm{Sub}_{U}[I_{\mathrm{no}}]$ are identically distributed, as desired. ∎
C.2 Concentration of Value for Random Instances

### v2

@@ 2.3 Irredundant Relations and Galois Duality
A fundamental result of Geiger [geiger1968closed] shows that if the notion of “relations generated by $\Gamma$” in Definition 2.1 is suitably relaxed, then the resulting set $\langle\Gamma\rangle^{*}$ actually coincides with $\langle\Gamma\rangle$. This result captures a deep “duality” between algebraic and relational structures (an instance of the more general concept of a Galois connection), which we do not elaborate on here. Instead, following Dalmau and Krokhin [dalmau2013robust], we state a variant of Geiger’s theorem asserting that any irredundant relation in $\langle\Gamma\rangle$ is contained in $\langle\Gamma\rangle^{*}$ (without relaxing Definition 2.1).
@@
Theorem 2.16 ([geiger1968closed]).
Let $(D,\Gamma)$ be a repetition-closed relational structure, and let $R:D^{k}\rightarrow\{0,1\}$ be an irredundant relation on $D$. If every polymorphism of $(D,\Gamma)$ preserves $R$, then $R\in\langle\Gamma\rangle^{*}$.
For the sake of completeness, a proof of Theorem 2.16 is given in Appendix C.
@@
For any repetition-closed relational structure $(D,\Gamma)$, we have $End_{\Gamma}\in\langle\Gamma\rangle^{*}$.
@@ 3 Main Lemma from Universal Algebra
The statement of Theorem 3.2 is primarily derived from [barto2014constraint, Conjecture 4.3], with additional ingredients drawn from other works, such as [valeriote2009subalgebra]. Further details concerning Theorem 3.2 will be provided in Appendix E.
@@
Let $(D,\Gamma)$ be a repetition-closed relational structure of unbounded width. Then there exist a subset $D^{\prime}\subseteq D$, an Abelian group $\mathsf{G}$, and a surjective map $\varphi\colon D^{\prime}\to\mathsf{G}$ such that any ternary relation $R\colon D^{3}\to\{0,1\}$ of the form
@@ 4 Hardness of Linear Equations
Theorem 4.1 ([bogdanov2002lower]).
For any Abelian group $\mathsf{G}$ and any $\varepsilon>0$, there exists a constant $d>0$ such that $\mathsf{MaxCSP}(\textsc{3Sum}_{\mathsf{G}})[1,\frac{1}{|\mathsf{G}|}+\varepsilon]$ requires $\Omega(n)$ queries in the $\mathrm{BD}^{*}(d,\frac{d}{3},3n)$ model. [^12] The paper [bogdanov2002lower] only proved Theorem 4.1 in the case $\mathsf{G}=\mathbb{Z}/2\mathbb{Z}$, but the proof can easily be adapted to work for general Abelian groups.
However, unlike [dalmau2013robust], which applies the result of [haastad2001some] as a black box to prove Theorem 1.9, we cannot directly use Theorem 4.1 due to the subtleties of our bounded-degree query model. In this section, we unpack the proof of Theorem 4.1 and extract the components that can be explicitly leveraged in proving Theorem 2.4. For the sake of convenience, several aspects of the proof in [bogdanov2002lower] are slightly modified in the presentation below.
@@
It is not hard to deduce Theorem 4.1 from Proposition 4.4 and Lemmas 4.5 and 4.7. For the sake of completeness, proofs of Lemmas 4.5 and 4.7 will be sketched in Appendix D.
@@ 5.1 The Expander Gadget
Recall that our universal-algebraic lemma (Lemma 3.4) guarantees the simulation of “linear equation constraints” only by $\Gamma\cup\textsc{Const}_{D}$, rather than by $\Gamma$ alone. Consequently, we must first find a way to simulate the constant relations using only the relations in $\Gamma$. Note that constant relations may not be generated by $\Gamma$ in the sense of Definition 2.1, so different techniques are required. Interestingly, the original work of [bogdanov2002lower] encountered the same difficulty when proving the hardness of testing graph 3-colorability. [^13] We now view this need for constant relations in [bogdanov2002lower] as a special case of the general phenomenon described in Lemma 3.4, since 3-colorability is an example of an unbounded-width CSP. To address this issue, [bogdanov2002lower] uses an “expander gadget” to fix the assignments of certain variables — a technique that also plays a crucial role in our proof.
In contrast to [bogdanov2002lower], where ordinary expander graphs suffice, our more general setting requires the use of random regular hypergraphs, defined formally below.
@@
A proof of Lemma 5.4 is given in Appendix D.
@@ 5.2 Construction in the Core Case
Recall that the expander gadget constructed in Section 5.1 relies on the availability of a sub-unique relation. However, some CSP templates cannot generate any sub-unique relation (in the sense of Definition 2.1). Fortunately, when the template $(D,\Gamma)$ is a core (see Definition 2.6), the “Endomorphism” relation provides a natural sub-unique relation that can be generated from $\Gamma$, as established in Proposition 2.18.151515The idea of using the Endomorphism relation is due to [dalmau2013robust].
In this subsection, we construct the YES and NO distributions for Theorem 2.4 in the case the template is a core. Throughout this subsection, we fix a repetition-closed relational structure $(D,\Gamma)$, and assume that it is a core and has unbounded width. In particular, this implies $|D|\geqslant 2$.
@@ 5.4 Conclusion in the General Case
We now turn to the case where $(D,\Gamma)$ is not a core. By Proposition 2.8, $(D,\Gamma)$ is homomorphically equivalent to a reptition-closed core $(D^{\prime},\Gamma^{\prime})$. We claim that $\mathsf{MaxCSP}(\Gamma^{\prime})[1,1-\varepsilon]$ is essentially equivalent to $\mathsf{MaxCSP}(\Gamma)[1,1-\varepsilon]$.
@@ 2   Initialize a set of partial assignments $S_{U}\leftarrow D^{U}$
Let $(D,\Gamma_{1})$ and $(D,\Gamma_{2})$ be two relational structures on a same domain. If $(D,\Gamma_{1})$ has unbounded width and $\Gamma_{1}\subseteq\Gamma_{2}$, then $(D,\Gamma_{2})$ also has unbounded width.
@@
This is simply because any $\mathrm{CSP}(\Gamma_{1})$ instance is also a $\mathrm{CSP}(\Gamma_{2})$ instance. If Algorithm 1 makes error on some instance $I\in\mathrm{CSP}(\Gamma_{1})$, then it also makes error on the same instance $I\in\mathrm{CSP}(\Gamma_{2})$. ∎
@@
A decidable characterization of bounded-width CSP templates was given by Barto [barto2014collapse, Corollary 8.5]. Using Barto’s criterion, it is easy to verify that for any $k,\ell\geqslant 2$ with $(k,\ell)\neq(2,2)$, the CSP template expressing $k$-colorability of $\ell$-uniform hypergraphs does not have bounded width.
Appendix B Allowing Variable Repetition
Proof of Theorem 1.8 assuming Theorem 2.4.
Given any relational structure $(D,\Gamma)$, we will define a reduction map $\mathfrak{R}:\mathrm{CSP}(\overline{\Gamma})\rightarrow\mathrm{CSP}(\Gamma)$. Let $k$ be the maximum arity of any relation in $\Gamma$. Given an instance $I=(V,\mathcal{C})\in\mathrm{CSP}(\overline{\Gamma})$, we construct $\mathfrak{R}[I]$ as follows.
  1. (1)
    Replace every variable $v\in V$ by $m:=k|D|$ copies $v^{(1)},\dots,v^{(m)}$. The new variable set is denoted by $V\times[m]$.
  2. (2)
    Each constraint $C\in\mathcal{C}$ can be viewed as a requirement
    $R(\tau(v_{1}),\dots,\tau(v_{r}))=1$
    for assignments $\tau:V\rightarrow D$, where $R\in\Gamma$ and $v_{1},\dots,v_{r}$ are not necessarily distinct variables in $V$. In place of this constraint, we add to $\mathfrak{R}[I]$ every constraint of the form
    $\left(\big(v_{1}^{(i_{1})},\dots,v_{r}^{(i_{r})}\big),R\right)$
    such that the chosen variables $v_{1}^{(i_{1})},\dots,v_{r}^{(i_{r})}$ are pairwise distinct.
The reduction map has the following desired properties.
  1. (1)
    Note that every constraint of $I$ corresponds to at least 1 and at most $m^{k}$ constraints of $\mathfrak{R}[I]$. If $I$ has maximum degree at most $d$, then $\mathfrak{R}[I]$ has maximum degree at most $m^{k}d$.
  2. (2)
    It is clear that if $I$ is satisfiable then $\mathfrak{R}[I]$ is also satisfiable.
  3. (3)
    If $\mathrm{val}_{I}\leqslant 1-\varepsilon$, then we claim that $\mathrm{val}_{\mathfrak{R}[I]}\leqslant 1-m^{-k}\varepsilon$. Given any assignment $\widetilde{\tau}:V\times[m]\rightarrow D$ and any $v\in V$, by the pigeonhole principle some element of $D$ appears in $\tau(v^{(1)}),\dots,\tau(v^{(m)})$ at least $k$ times. We define $\tau(v)$ to be any such element. For any constraint $C\in\mathcal{C}$ that this assignment $\tau:V\rightarrow D$ violates, at least one of the (at most $m^{k}$) constraints in $\mathfrak{R}[I]$ corresponding to $C$ is violated by $\widetilde{\tau}$. Therefore, the assumption that $\mathrm{val}_{I}(\tau)\leqslant 1-\varepsilon$ implies $\mathrm{val}_{\mathfrak{R}[I]}(\widetilde{\tau})\leqslant 1-m^{-k}\varepsilon$.
  4. (4)
    Finally, it is easy to see that every query to a variable $v^{(i)}$ in $\mathfrak{R}[I]$ reveals no more information than a query to $v$ in $I$.
Now suppose Theorem 2.4 holds. Then for any unbounded-width CSP template $(D,\Gamma)$, since $(D,\overline{\Gamma})$ is repetition-closed and also has unbounded width (due to Proposition A.4), there exists $\varepsilon>0$ such that $\mathsf{MaxCSP}(\overline{\Gamma})[1,1-\varepsilon]$ requires $\Omega(n)$ query in the $\mathrm{BD}^{*}(d,\alpha,n)$ model. Due to the reduction above, this implies that $\mathsf{MaxCSP}(\Gamma)[1,1-m^{-k}\varepsilon]$ requires $\Omega(n)$ queries in the $\mathrm{BD}^{*}(m^{k}d,\alpha,mn)$ model, where $k$ is the largest arity of any relation in $\Gamma$ and $m:=k|D|$. ∎
Appendix C Galois Duality Proof
@@
Appendix D Random Regular Hypergraphs
In this section, we outline the proofs of Lemmas 4.5, 4.7 and 5.4, all of which revolve around random regular hypergraphs. The first two lemmas, Lemmas 4.5 and 4.7, are only minor adaptations of results from [bogdanov2002lower], and our proofs closely follow those in that work. In contrast, Lemma 5.4 is new to this paper, though its proof uses arguments that are largely analogous to those used in the proof of Lemma 4.5.
D.1 Random Regular Hypergraphs are Locally Sparse
@@
Recall from Definition 4.2 that $\mathcal{M}_{[3],n}$ denotes the set of all $3$-partite perfect matchings on the vertex set $n]\times[3]$. For any matching $M\in\mathcal{M}_{[3],n}$ and any vertex subset $U\subseteq[n]\times[3]$, we let $M[U]$ denote the set of hyperedges $(v_{1},v_{2},v_{3})\in M$ such that $v_{1},v_{2},v_{3}\in U$. The following lemma is a slightly modified version of [bogdanov2002lower, Lemma 6].
Lemma D.1 ([bogdanov2002lower]).
@@
$\sum_{i=1}^{d}\left|M^{(i)}[U]\right|<\frac{2}{3}|U|,\quad\text{for all }U\subseteq[n]\times[3]\text{ with }|U|\leqslant\delta n$ (D.1)
@@
As stated in [bogdanov2002lower], the conclusion of Lemma D.1 still holds if the constant $2/3$ in (D.1) is replaced by any constant greater than $1/2$. The reason for choosing $2/3$ in (D.1) is that it is the largest constant for which the following observation remains true.
Proposition D.2.
Suppose $M^{(1)},\dots,M^{(d)}$ are perfect matchings in $\mathcal{M}_{[3],n}$ that satisfy (D.1). Then, for any subset $U\subseteq[n]\times[3]$ with $|U|\leqslant\delta n$, the following hold:
@@
The first item holds because no hyperedge can appear in more than one of the $M^{(i)}$’s; otherwise, there would exist a set $U$ of three vertices such that $\sum_{i=1}^{d}\left|M^{(i)}[U]\right|\geqslant 2$, violating condition (D.1).
We now prove item (2). Let $U_{0}\subseteq U$ denote the set of vertices incident to at least one hyperedge in the union $E:=\bigcup_{i=1}^{d}M^{(i)}$. Applying condition (D.1) to $U_{0}$ implies that a uniformly random vertex from $U_{0}$ is, in expectation, incident to fewer than two hyperedges in $E$. Hence, there exists at least one vertex in $U_{0}$ that is incident to exactly one hyperedge in $E$.
@@
We now use Lemma D.1 and Proposition D.2 to prove Lemma 4.7.
@@
Recall from Definition 4.3 that for any sample $(I_{\mathrm{yes}},I_{\mathrm{no}})\sim\mathcal{D}_{\mathrm{pair}}^{\mathsf{G},n,d}$, the same collection of $d$ perfect matchings $M^{(1)},\dots,M^{(d)}\in\mathcal{M}_{[3],n}$ underlies both $I_{\mathrm{yes}}$ and $I_{\mathrm{no}}$. Applying Lemma D.1, we obtain a constant $\delta\in(0,1)$, and let $\mathcal{E}$ denote the event that the condition (D.1) holds for $M^{(1)},\dots,M^{(d)}$.
We now condition on a choice of $M^{(1)},\dots,M^{(d)}$ satisfying (D.1). Recall that for each hyperedge $\mathbf{v}=(v_{1},v_{2},v_{3})\in\bigcup_{i=1}^{d}M^{(i)}$, the instance $I_{\mathrm{yes}}$ contains a constraint $(\mathbf{v},S^{\mathsf{G}}_{b})$ and $I_{\mathrm{no}}$ contains a constraint $(\mathbf{v},S^{\mathsf{G}}_{b^{\prime}})$, for possibly distinct “right-hand side” coefficients $b,b^{\prime}\in\mathsf{G}$. The right-hand sides of the equations in $I_{\mathrm{no}}$ are chosen independently, whereas those in $I_{\mathrm{yes}}$ are correlated through an underlying random assignment $\tau:[n]\times[3]\to\mathsf{G}$.
However, it follows from Proposition D.2 that for any variable set $U\subseteq[n]\times[3]$ of size at most $\delta n$, the equations placed by $I_{\mathrm{yes}}$ within $U$ have right-hand sides that remain jointly independent. Consequently, conditioned on $M^{(1)},\dots,M^{(d)}$, the restricted random instances $\mathrm{Sub}_{U}[I_{\mathrm{yes}}]$ and $\mathrm{Sub}_{U}[I_{\mathrm{no}}]$ are identically distributed, as desired. ∎
D.2 Concentration of Value for Random Instances